# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 1 Introduction
Such turbulent heating prescriptions are also critical for the interpretation of remote astronomical observations of emissions from black hole accretion disks, such as the ground-breaking observations by the Event Horizon Telescope of the supermassive black holes at the center of M87 (Event Horizon Telescope Collaboration et al., 2019) and at Sagittarius $*{}^{*}$ in the Milky Way (Event Horizon Telescope Collaboration et al., 2022), In these cases, alternative turbulent heating prescriptions (Howes, 2010; Rowan et al., 2017) have been shown to yield drastically different predictions for the emitted radiation (Chael et al., 2019).
@@ 1.1 How Does Turbulence Mediate Energy Transport and Plasma Heating?
The transport of energy mediated by plasma turbulence is depicted in Fig. 1, where (a) a plot of the turbulent magnetic energy wavenumber spectrum shows that energy injected into the turbulent cascade at the outer scale ($k_{0}\rho_{i}=10^{-4}$) is transferred nonlinearly without loss through the inertial range to the transition into the dissipation range at ion scales ($k\rho_{i}\sim 1$). At the ion scales and smaller ($k\rho_{i}\gtrsim 1$), poorly understood mechanisms can remove energy from the turbulence, transferring energy to the ions (blue arrows) or electrons (magenta arrows). In Fig. 1(b), modern plasma turbulence theory [(Goldreich & Sridhar, 1995, hereafter GS95) or (Boldyrev, 2006, hereafter B06)] predicts that the turbulent cascade generates a scale-dependent anisotropy in wavevector space (blue shaded region); isotropy at the driving scale ($k_{\perp 0}\rho_{i}=k_{\parallel 0}\rho_{i}=10^{-4}$) develops into a significant anisotropy $k_{\perp}\gg k_{\parallel}$ as the cascade transfers energy to smaller scales [^1] Parallel and perpendicular here are defined relative to the local mean magnetic field direction.. In addition, kinetic instabilities can generate unstable turbulent fluctuations with wavevectors $k\rho_{i}\sim 1$ (red ovals in both panels), which can mediate the nonlocal transfer of energy from the large-scale motions directly to the ion kinetic scales. I explicitly define “turbulence” here as the collection of physical mechanisms involved in the conversion of the energy of large-scale plasma flows and magnetic fields into plasma heat or the energy of accelerated particles, including the nonlinear scale-to-scale energy transfer of the turbulent cascade, the mechanisms damping the turbulence at small scales, and any possible instabilities that may generate turbulent fluctuations or govern nonlocal energy transfer.
@@ 1.2 What Do We Want to Predict?
In a statistically steady-state turbulent cascade in a weakly collisional astrophysical plasma, the energy injected into the turbulence at the outer scale is transferred locally in scale down to increasingly smaller scales until reaching the kinetic length scales at which damping mechanisms can collisonlessly remove energy from the turbulence and energize the different particle species, as depicted in Fig. 1(a). This local cascade of energy is mediated by nonlinear interactions among the turbulent fluctuations, such as between counterpropagating Alfvén wave packets, often denoted “Alfvén wave collisions” (Kraichnan, 1965; Sridhar & Goldreich, 1994; Goldreich & Sridhar, 1995; Ng & Bhattacharjee, 1996; Galtier et al., 2000; Howes et al., 2012b; Howes & Nielson, 2013; Nielson et al., 2013; Howes et al., 2013; Drake et al., 2013; Howes, 2016; Verniero et al., 2018; Verniero & Howes, 2018; TenBarge et al., 2021; Ripperda et al., 2021). Nonlocal energy transfer mechanisms—such as kinetic instabilities driven by the large-scale dynamics that can cause the plasma conditions to deviate from local thermodynamic equilibrium—can also directly generate turbulent fluctuations at the small kinetic length scales, as illustrated in Fig. 1(b). Together, these mechanisms lead to a turbulent energy density cascade rate $\varepsilon$ from the large scales down to the small scales of the turbulence. This nonlinear transfer of the turbulent energy density is ultimately damped at small scales by kinetic plasma physics mechanisms, converting the energy density of the small-scale turbulent plasma flows and magnetic fields into particle energy with a statistically steady-state total plasma energy density heating rate $Q\sim\varepsilon$.
@@
It is worthwhile noting that the determination of the partitioning of the turbulent energy between ions and electrons $Q_{i}/Q_{e}$ can be simplified by the properties of the turbulent cascade that are evident in Fig. 1(a). The inertial range is defined as the range of scales over which the turbulent energy density is transferred to smaller scales without loss, leading to a constant energy density cascade rate $\varepsilon$ throughout this range [^3] This standard definition of the inertial range from fluid turbulence theory neglects the possibility of a nonlocal transfer of energy that may arise in weakly collisional plasmas.. Therefore, the turbulent cascade rate $\varepsilon$ at the onset of damping at ion kinetic scales at $k\rho_{i}\sim 1$ is the same as that determined at the outer scale of the turbulence at $k_{0}\rho_{i}\ll 1$. The collisionless wave-particle interactions that serve to remove energy from the turbulent fluctuations and transfer that energy to the ions generally play a significant role only within the range of length scales similar to the ion kinetic length scales, $k\rho_{i}\sim 1$: (i) at larger scales $k\rho_{i}\ll 1$, the collisionless damping rate with the ions $\gamma_{i}$ is typically negligible in comparison to the turbulent cascade rate $\gamma_{i}\ll\omega_{nl}$; and (ii) at smaller scales $k\rho_{i}\gg 1$, the gyromotion of the ions averages out the effect of the small-scale electromagnetic fluctuations, so the ions decouple from the turbulent fields and do not exchange significant energy. Whatever turbulent energy manages to cascade beyond the ion scales to smaller scales at $k\rho_{i}\gg 1$ will ultimately be deposited with the electrons. Thus, the task of determining $Q_{i}/Q_{e}$ boils down to determining what fraction of the turbulent energy is damped on the ions as the cascade progresses through the ion kinetic scales at $k\rho_{i}\sim 1$.
@@ 2 The Fundamental Dimensionless Parameters of Kinetic Plasma Turbulence
The first step in developing a predictive model for plasma turbulence in weakly collisional space and astrophysical plasmas is to identify the key dimensionless parameters upon which the turbulent energy cascade and its dissipation depend. In this paper, the focus is to characterize kinetic plasma turbulence in the non-relativistic limit. There are two categories of governing parameters: Plasma Parameters and Turbulence Parameters. Modern scaling theories for anisotropic plasma turbulence can be employed to reduce the number of parameters needed to characterize the turbulence, as described in §3.1. The reduced set of dimensionless parameters can then be used to develop useful phase diagrams that identify the dominant physical mechanisms of turbulent dissipation and which of these key parameters has the strongest impact on how particles are energized by the turbulence.
@@ 2.1 Plasma Parameters
Although for a hydrogenic plasma of protons and electrons the ion-to-electron mass ratio $\mu=m_{i}/m_{e}$ has a fixed physical value of 1836, this ratio is included as an additional independent parameter in the calculations below since a reduced mass ratio is often used in numerical simulations of plasma turbulence for reasons of computational efficiency. Note also that, once $\mu$ and $\tau$ are specified, the single dimensionless parameter $k_{\parallel 0}\lambda_{mfp,e}$ is sufficient to specify all possible charged particle collision rates, $\nu_{ee}$, $\nu_{ei}$, $\nu_{ii}$, and $\nu_{ie}$.
@@
Parallel Ion Plasma Beta $\beta_{\parallel,i}=8\pi n_{i}T_{\parallel i}/B^{2}$
@@ 2.2 Turbulence Parameters
The Turbulence Parameters characterize the scale, amplitude, and properties of the turbulent driving though five additional parameters. For simplicity, the focus here is restricted to only the case of turbulence in a spatially homogeneous system in which the scale length of gradients in the equilibrium quantities are much larger than the outer scale of the system. Dimensionless ratios of length scales are notationally simplified by characterizing the length-scales of the turbulent fluctuations in terms of their perpendicular and parallel wavenumbers. Below quantities at the outer scale of the turbulence, commonly denoted the driving scale or energy injection scale, are denoted by the subscript “0”.
@@
The driving of the turbulence can be characterized by three dimensionless parameters. The first parameter is the perpendicular wavenumber of the driving-scale fluctuations normalized by the ion Larmor radius, $k_{\perp 0}\rho_{i}$. The second parameter is the wavevector anisotropy of the fluctuations at the driving scale, $k_{\parallel 0}/k_{\perp 0}$. The third parameter is the the nonlinearity parameter that characterizes the amplitude of the driving, $\chi_{0}=(k_{\perp 0}\delta B_{\perp 0}\theta_{0})/(k_{\parallel 0}B_{0})$, where $\delta B_{\perp 0}$ is the amplitude of the perpendicular magnetic field fluctuations, $B_{0}$ is the equilibrium magnetic field magnitude, and $\theta_{0}$ is the angle of dynamic alignment at the driving scale in radians (Boldyrev, 2006). In the next section, we demonstrate how modern scaling theories for anisotropic MHD turbulence can be used to combine these three dimensionless parameters into a single parameter, the isotropic driving wavenumber, $k_{0}\rho_{i}$. Note also that, in principle, the dynamic alignment angle $\theta_{0}$ at driving scale is an additional dimensionless parameter, but dynamic alignment is typically understood to be a phenomenon that develops self-consistently due to the turbulence as it cascades to smaller scales, so we neglect this parameter at the outerscale, generally taking $\theta_{0}\sim 1$.
@@
The nature of the turbulent fluctuations at the driving scale also plays an important role in determining the ultimate fate of the turbulent energy. The fourth turbulence parameter is the imbalance parameter, measuring the ratio of the amplitudes in upward-to-downward Elsasser fields, $Z_{0}^{+}/Z_{0}^{-}$, where the Elsasser fields at the outer scale are defined by $Z_{0}^{\pm}=\delta\mathbf{U}_{\perp,0}\pm\delta B_{\perp,0}/\sqrt{4\pi n_{i}m_{i}}$ (Elsasser, 1950), and where $\delta\mathbf{U}_{\perp,0}$ and $\delta B_{\perp,0}$ are the perpendicular perturbed plasma flow velocity and perpendicular perturbed magnetic field of the turblent fluctuations at the driving scale. Physically, Alfvén wave packets propagating up the mean magnetic field have $Z^{+}=0$ and $Z^{-}\neq 0$, and those propagating down the field have $Z^{+}\neq 0$ and $Z^{-}=0$. The Elsasser fields effectively characterize the wave energy flux in each direction along the magnetic field [^5] Note that the imbalance of counterpropagating wave energy fluxes in plasma turbulence has alternatively been characterized in the literature using the cross helicity (Dobrowolny et al., 1980; Marsch & Tu, 1990; Matthaeus et al., 1982; Perez & Boldyrev, 2009, 2010) or the normalized Poynting flux (Hnat et al., 2003; Di Mare & Howes, 2024). of the typically dominant Alfvénic fluctuations in heliospheric plasma turbulence (Tu & Marsch, 1995; Schekochihin et al., 2009; Bruno & Carbone, 2013). Since the nonlinear interactions that mediate the turbulent cascade in Alfvénic turbulence at MHD scales arise only between counterpropagating and perpendicularly polarized Alfvén waves (Kraichnan, 1965; Sridhar & Goldreich, 1994; Goldreich & Sridhar, 1995; Ng & Bhattacharjee, 1996; Galtier et al., 2000; Howes & Nielson, 2013; Howes:2014c), the imbalance plays an important role in characterizing the turbulence. The importance of $Z_{0}^{+}/Z_{0}^{-}$ under sufficiently imbalanced conditions, such as occurs in the inner heliosphere where wave energy fluxes are dominantly anti-sunward, has been demonstrated clearly with the recent discovery of new phenomenon known as the helicity barrier in plasma turbulence (Squire et al., 2022, 2023).
@@
A fifth turbulence parameter is the ratio of the energy in compressible fluctuations to that in incompressible fluctuations, $E_{comp}/E_{inc}$. Since a magnetized plasma supports two distinct compressible wave modes, the fast and slow magnetosonic waves (differing by whether the thermal and magnetic pressure perturbations associated with the wave act in conjunction or in opposition), one could further characterize the compressible component of the turbulent fluctuations by the additional ratio of the energy in the fast magnetosonic fluctuations to the total energy of compressible fluctuations $E_{F}/E_{comp}$, where the total compressible energy is the sum of the contrbutions from fast-mode and slow-mode fluctuations, $E_{comp}=E_{F}+E_{s}$. The ratio $E_{F}/E_{comp}$ is not included here as one of the fundamental parameters based on a study of the inertial range in solar wind turbulence at 1 AU that showed negligible energy in fast-mode mode fluctuations, or $E_{F}/E_{comp}\ll 1$ (Howes et al., 2012a). In significant energy exists in fast-mode fluctuations in other turbulent astrophysical plasmas, $E_{F}/E_{comp}$ may need to be included.

### v2

@@ 1 Introduction
Such turbulent heating prescriptions are also critical for the interpretation of remote astronomical observations of emissions from black hole accretion disks, such as the ground-breaking observations by the Event Horizon Telescope of the supermassive black holes at the center of M87 (Event Horizon Telescope Collaboration et al., 2019) and at Sagittarius A in the Milky Way (Event Horizon Telescope Collaboration et al., 2022). In these cases, alternative turbulent heating prescriptions (Howes, 2010; Rowan et al., 2017) have been shown to yield drastically different predictions for the emitted radiation (Chael et al., 2019).
@@ 1.1 How Does Turbulence Mediate Energy Transport and Plasma Heating?
The transport of energy mediated by plasma turbulence is depicted in Fig. 1, where (a) a plot of the turbulent magnetic energy wavenumber spectrum shows that energy injected into the turbulent cascade at the outer scale ($k_{0}\rho_{i}=10^{-4}$) is transferred nonlinearly without loss through the inertial range to the transition into the dissipation range at ion scales ($k\rho_{i}\sim 1$). At the ion scales and smaller ($k\rho_{i}\gtrsim 1$), poorly understood mechanisms can remove energy from the turbulence, transferring energy to the ions (blue arrows) or electrons (magenta arrows). In Fig. 1(b), modern plasma turbulence theory [(Goldreich & Sridhar, 1995, hereafter GS95) or (Boldyrev, 2006, hereafter B06)] predicts that the turbulent cascade generates a scale-dependent anisotropy in wavevector space (blue shaded region): isotropy at the driving scale ($k_{\perp 0}\rho_{i}=k_{\parallel 0}\rho_{i}=10^{-4}$) develops into a significant anisotropy $k_{\perp}\gg k_{\parallel}$ as the cascade transfers energy to smaller scales [^1] Parallel and perpendicular here are defined relative to the local mean magnetic field direction.. In addition, kinetic instabilities can generate unstable turbulent fluctuations with wavevectors $k\rho_{i}\sim 1$ (red ovals in both panels), which can mediate the nonlocal transfer of energy from the large-scale motions directly to the ion kinetic scales. I explicitly define “turbulence” here as the collection of physical mechanisms involved in the conversion of the energy of large-scale plasma flows and magnetic fields into plasma heat or the energy of accelerated particles, including the nonlinear scale-to-scale energy transfer of the turbulent cascade, the mechanisms damping the turbulence at small scales, and any possible instabilities that may generate turbulent fluctuations or govern nonlocal energy transfer.
@@ 1.2 What Do We Want to Predict?
In a statistically steady-state turbulent cascade in a weakly collisional astrophysical plasma, the energy injected into the turbulence at the outer scale is transferred locally in scale down to increasingly smaller scales until reaching the kinetic length scales at which damping mechanisms can collisionlessly remove energy from the turbulence and energize the different particle species, as depicted in Fig. 1(a). This local cascade of energy is mediated by nonlinear interactions among the turbulent fluctuations, such as between counterpropagating Alfvén wave packets, often denoted “Alfvén wave collisions” (Kraichnan, 1965; Sridhar & Goldreich, 1994; Goldreich & Sridhar, 1995; Ng & Bhattacharjee, 1996; Galtier et al., 2000; Howes et al., 2012b; Howes & Nielson, 2013; Nielson et al., 2013; Howes et al., 2013; Drake et al., 2013; Howes, 2016; Verniero et al., 2018; Verniero & Howes, 2018; TenBarge et al., 2021; Ripperda et al., 2021). Nonlocal energy transfer mechanisms—such as kinetic instabilities driven by the large-scale dynamics that may cause the plasma conditions to deviate from local thermodynamic equilibrium—can also directly generate turbulent fluctuations at the small kinetic length scales, as illustrated in Fig. 1(b). Together, these mechanisms lead to a turbulent energy density cascade rate $\varepsilon$ from the large scales down to the small scales of the turbulence. This nonlinear transfer of the turbulent energy density is ultimately damped at small scales by kinetic plasma physics mechanisms, converting the energy density of the small-scale turbulent plasma flows and magnetic fields into particle energy with a statistically steady-state total plasma energy density heating rate $Q\sim\varepsilon$.
@@
It is worthwhile noting that the determination of the partitioning of the turbulent energy between ions and electrons $Q_{i}/Q_{e}$ can be simplified by the properties of the turbulent cascade that are evident in Fig. 1(a). The inertial range is defined as the range of scales over which the turbulent energy density is transferred to smaller scales without loss, leading to a constant energy density cascade rate $\varepsilon$ throughout this range [^3] This standard definition of the inertial range from fluid turbulence theory neglects the possibility of a nonlocal transfer of energy that may arise in weakly collisional plasmas.. Therefore, the turbulent cascade rate $\varepsilon$ at the onset of damping at ion kinetic scales at $k\rho_{i}\sim 1$ is the same as that determined at the outer scale of the turbulence at $k_{0}\rho_{i}\ll 1$. The collisionless wave-particle interactions that serve to remove energy from the turbulent fluctuations and transfer that energy to the ions generally play a significant role only within the range of length scales similar to the ion kinetic length scales, $k\rho_{i}\sim 1$: (i) at larger scales $k\rho_{i}\ll 1$, the collisionless damping rate with the ions $\gamma_{i}$ is typically negligible in comparison to the effective nonlinear frequency of the turbulent energy transfer $\gamma_{i}\ll\omega_{nl}$; and (ii) at smaller scales $k\rho_{i}\gg 1$, the gyromotion of the ions averages out the effect of the small-scale electromagnetic fluctuations, so the ions decouple from the turbulent fields and do not exchange significant energy. Whatever turbulent energy manages to cascade beyond the ion scales to smaller scales at $k\rho_{i}\gg 1$ will ultimately be deposited with the electrons. Thus, the task of determining $Q_{i}/Q_{e}$ boils down to determining what fraction of the turbulent energy is damped on the ions as the cascade progresses through the ion kinetic scales at $k\rho_{i}\sim 1$.
@@ 2 The Fundamental Dimensionless Parameters of Kinetic Plasma Turbulence
The first step in developing a predictive model for plasma turbulence in weakly collisional space and astrophysical plasmas is to identify the key dimensionless parameters upon which the turbulent energy cascade and its dissipation depend. In this paper, the focus is to characterize kinetic plasma turbulence in the non-relativistic limit for sub-Alfvénic turbulent motions, a case widely applicable to the turbulent plasmas found ubiquitously in the heliosphere and throughout the universe. There are two categories of governing parameters: Plasma Parameters and Turbulence Parameters. Modern scaling theories for anisotropic plasma turbulence can be employed to reduce the number of parameters needed to characterize the turbulence, as described in §3.1. The reduced set of dimensionless parameters can then be used to develop useful phase diagrams that identify the dominant physical mechanisms of turbulent dissipation and highlight which of these key parameters has the strongest impact on how particles are energized by the turbulence.
Characteristic Plasma Quantities
Parameter Symbol and Definition
Parallel Species Thermal Velocity $v_{t\parallel,s}=\sqrt{2T_{\parallel,s}/m_{s}}$
Perpendicular Species Thermal Velocity $v_{t\perp,s}=\sqrt{2T_{\perp,s}/m_{s}}$
Alfvén Velocity $v_{A}=B_{0}/\sqrt{4\pi n_{i}m_{i}}$
Ion Acoustic Velocity $c_{s}=\sqrt{T_{\parallel,e}/m_{i}}$
Species Cyclotron Frequency (Angular) $\Omega_{s}=q_{s}B_{0}/m_{s}c$
Species Plasma Frequency (Angular) $\omega_{ps}=\sqrt{4\pi n_{s}q_{s}^{2}/m_{s}}$
Electron-electron Collision Frequency $\nu_{ee}=2^{3/2}\pi n_{e}e^{4}\ln\Lambda/(m_{e}^{1/2}T_{\parallel,e}^{3/2})$
Ion Larmor Radius $\rho_{i}=v_{t\perp,i}/\Omega_{i}$
Ion Sound Larmor Radius $\rho_{s}=c_{s}/\Omega_{i}$
Electron Larmor Radius $\rho_{e}=v_{t\perp,e}/\Omega_{e}$
Species Inertial Length $d_{s}=c/\omega_{ps}$
Species Debye Length $\lambda_{D,s}=\sqrt{T_{s}/(4\pi n_{s}q_{s}^{2})}$
Electron Mean Free Path $\lambda_{mfp,e}=v_{t\parallel,e}/\nu_{ee}$
Ion Mean Free Path $\lambda_{mfp,i}=v_{t\parallel,i}/\nu_{ii}$
Plasma Parameter $\Lambda=n_{e}\lambda_{D,e}^{3}$
Parallel Driving Scale $L_{\parallel 0}$
Perpendicular Driving Scale $L_{\perp 0}$
Parallel Driving Wavenumber $k_{\parallel 0}=2\pi/L_{\parallel 0}$
Perpendicular Driving Wavenumber $k_{\perp 0}=2\pi/L_{\perp 0}$
Useful Conversion Formulas
$d_{i}=c/\omega_{pi}=v_{A}/\Omega_{i}=\rho_{i}/\sqrt{\beta_{\parallel,i}A_{i}}$ $\nu_{ei}\sim\nu_{ee}$
$d_{e}=v_{A}\mu^{1/2}/\Omega_{e}=\rho_{i}/\sqrt{\beta_{\parallel,i}A_{i}\mu}=\rho_{e}\sqrt{\tau_{\parallel}\mu/(\beta_{\parallel,i}A_{e})}$ $\nu_{ii}\sim\mu^{-1/2}\tau_{\parallel}^{-3/2}\nu_{ee}$
$\rho_{s}=\rho_{i}/\sqrt{2\tau_{\parallel}A_{i}}$ $\nu_{ie}\sim\mu^{-1}\tau_{\parallel}^{-3/2}\nu_{ee}$
$\rho_{e}=\rho_{i}/\sqrt{\tau_{\parallel}A_{e}\mu/A_{i}}$ $\lambda_{mfp,i}=\lambda_{mfp,e}\tau_{\parallel}^{2}$
$\beta_{\parallel,i}=v_{t\parallel,i}^{2}/v_{A}^{2}$ $\beta_{\parallel,e}=\beta_{\parallel,i}/\tau_{\parallel}$
Table 2: Definitions of characteristic plasma quantities in cgs units and useful conversion relations, where $\beta_{\parallel,i}$, $\tau_{\parallel}$, $A_{i}$, and $\mu$ are defined in Table 3.
@@ 2.1 Plasma Parameters
Although for a hydrogenic plasma of protons and electrons the ion-to-electron mass ratio $\mu=m_{i}/m_{e}$ has a fixed physical value of 1836, this ratio is included as an additional independent parameter in the calculations below since a reduced mass ratio is often used in numerical simulations of plasma turbulence for reasons of computational efficiency. Note also that, once $\mu$ and $\tau_{\parallel}$ are specified, the single dimensionless parameter $k_{\parallel 0}\lambda_{mfp,e}$ is sufficient to specify all possible charged particle collision rates, since $\nu_{ee}/(k_{\parallel 0}v_{t\parallel,e})=1/(k_{\parallel 0}\lambda_{mfp,e})$ and $\nu_{ei}\sim\nu_{ee}$, $\nu_{ii}\sim\mu^{-1/2}\tau_{\parallel}^{-3/2}\nu_{ee}$, and $\nu_{ie}\sim\mu^{-1}\tau_{\parallel}^{-3/2}\nu_{ee}$.
@@
Parallel Ion Plasma Beta $\beta_{\parallel,i}=8\pi n_{i}T_{\parallel,i}/B^{2}$
@@ 2.2 Turbulence Parameters
The Turbulence Parameters characterize the scale, amplitude, and properties of the turbulent driving though five additional parameters. For simplicity, the focus here is restricted to only the case of turbulence in a spatially homogeneous system in which the scale length of gradients in the equilibrium quantities are much larger than the turbulent outer scale of the system. Dimensionless ratios of length scales are notationally simplified by characterizing the length-scales of the turbulent fluctuations in terms of their perpendicular and parallel wavenumbers. Below quantities at the outer scale of the turbulence, commonly denoted the driving scale or energy injection scale, are denoted by the subscript “0”.
@@
The driving of the turbulence can be characterized by three dimensionless parameters. The first parameter is the perpendicular wavenumber of the driving-scale fluctuations normalized by the ion Larmor radius, $k_{\perp 0}\rho_{i}$. The second parameter is the wavevector anisotropy of the fluctuations at the driving scale, $k_{\parallel 0}/k_{\perp 0}$. The third parameter is the nonlinearity parameter that characterizes the amplitude of the driving, $\chi_{0}=(k_{\perp 0}\delta B_{\perp 0}\theta_{0})/(k_{\parallel 0}B_{0})$, where $\delta B_{\perp 0}$ is the amplitude of the perpendicular magnetic field fluctuations, $B_{0}$ is the equilibrium magnetic field magnitude, and $\theta_{0}$ is the angle of dynamic alignment at the driving scale in radians (Boldyrev, 2006). In the next section, we demonstrate how modern scaling theories for anisotropic MHD turbulence can be used to combine these three dimensionless parameters into a single parameter, the isotropic driving wavenumber, $k_{0}\rho_{i}$. Note also that, in principle, the dynamic alignment angle $\theta_{0}$ at driving scale is an additional dimensionless parameter (measured in radians), but dynamic alignment is typically understood to be a phenomenon that develops self-consistently due to the turbulence as it cascades to smaller scales, so we neglect this parameter at the outer scale, generally taking $\theta_{0}\sim 1$ rad.
@@
The nature of the turbulent fluctuations at the driving scale also plays an important role in determining the ultimate fate of the turbulent energy. The fourth turbulence parameter is the imbalance parameter, measuring the ratio of the amplitudes in upward-to-downward Elsasser fields, $Z_{0}^{+}/Z_{0}^{-}$, where the Elsasser fields at the outer scale are defined by $\mathbf{Z}_{0}^{\pm}=\delta\mathbf{U}_{\perp,0}\pm\delta\mathbf{B}_{\perp,0}/\sqrt{4\pi n_{i}m_{i}}$ (Elsasser, 1950), and where $\delta\mathbf{U}_{\perp,0}$ and $\delta\mathbf{B}_{\perp,0}$ are the perpendicular perturbed plasma flow velocity and perpendicular perturbed magnetic field of the turbulent fluctuations at the driving scale. Formally, $\delta\mathbf{U}_{\perp,0}$ and $\delta\mathbf{B}_{\perp,0}$ are vectors in the plane perpendicular to the equilibrium magnetic field $\mathbf{B}_{0}$, but to calculate the ratio $Z_{0}^{+}/Z_{0}^{-}$, we simply take the amplitudes of the Elsasser fields, $Z_{0}^{\pm}=|\mathbf{Z}_{0}^{\pm}|$. Physically, Alfvén wave packets propagating up the mean magnetic field have $Z^{+}=0$ and $Z^{-}\neq 0$, and those propagating down the field have $Z^{+}\neq 0$ and $Z^{-}=0$. The Elsasser fields effectively characterize the wave energy flux in each direction along the magnetic field [^5] Note that the imbalance of counterpropagating wave energy fluxes in plasma turbulence has alternatively been characterized in the literature using the cross helicity (Dobrowolny et al., 1980; Marsch & Tu, 1990; Matthaeus et al., 1982; Perez & Boldyrev, 2009, 2010) or the normalized Poynting flux (Hnat et al., 2003; Di Mare & Howes, 2024). of the typically dominant Alfvénic fluctuations in heliospheric plasma turbulence (Tu & Marsch, 1995; Schekochihin et al., 2009; Bruno & Carbone, 2013). Since the nonlinear interactions that mediate the turbulent cascade in Alfvénic turbulence at MHD scales arise only between counterpropagating and perpendicularly polarized Alfvén waves (Kraichnan, 1965; Sridhar & Goldreich, 1994; Goldreich & Sridhar, 1995; Ng & Bhattacharjee, 1996; Galtier et al., 2000; Howes & Nielson, 2013; Howes, 2015), the imbalance plays an important role in characterizing the turbulence. The importance of $Z_{0}^{+}/Z_{0}^{-}$ under sufficiently imbalanced conditions, such as occurs in the inner heliosphere where wave energy fluxes are dominantly anti-sunward, has been demonstrated clearly with the recent discovery of a new phenomenon known as the helicity barrier in plasma turbulence (Meyrand et al., 2021; Squire et al., 2022, 2023).
@@
A fifth turbulence parameter is the ratio of the energy in compressible fluctuations to that in incompressible fluctuations, $E_{comp}/E_{inc}$. Since a magnetized plasma supports two distinct compressible wave modes, the fast and slow magnetosonic waves (differing by whether the thermal and magnetic pressure perturbations associated with the wave act in conjunction or in opposition), one could further characterize the compressible component of the turbulent fluctuations by the additional ratio of the energy in the fast magnetosonic fluctuations to the total energy of compressible fluctuations $E_{F}/E_{comp}$, where the total compressible energy is the sum of the contributions from fast-mode and slow-mode fluctuations, $E_{comp}=E_{F}+E_{s}$. The ratio $E_{F}/E_{comp}$ is not included here as one of the fundamental parameters based on a study of the inertial range in solar wind turbulence at 1 AU that showed negligible energy in fast-mode mode fluctuations, or $E_{F}/E_{comp}\ll 1$ (Howes et al., 2012a). In significant energy exists in fast-mode fluctuations in other turbulent astrophysical plasmas, $E_{F}/E_{comp}$ may need to be included.
@@
2.3 What Happened to the Reynolds Number?
Readers familiar with studies of MHD turbulence may find what appear to be two glaring omissions in the proposed list of Plasma and Turbulence parameters in Table 3 since it does not include the Reynolds number, $\mbox{Re}=LU_{0}/\nu$, or the magnetic Reynolds number, $\mbox{Re}_{M}=LU_{0}/\eta$. The Reynolds number is a dimensionless quantity that characterizes the ratio of the amplitude of the convective term $\mathbf{U}\cdot\nabla\mathbf{U}$ to that of the viscous diffusion term $\nu\nabla^{2}\mathbf{U}$ in the MHD momentum equation; the magnetic Reynolds number characterizes the ratio of the amplitude of the inductive term ×$\nabla\times(\mathbf{U}\times\mathbf{B})$ to that of the resistive diffusion term $\eta\nabla^{2}\mathbf{B}$ in the MHD induction equation.
The kinematic viscosity $\nu$ and resistivity $\eta$ are transport coefficients that are rigourously defined by an extension of the Chapman-Enskog procedure for magnetohydrodynamic systems (Spitzer, 1962; Grad, 1963; Braginskii, 1965) in the limit of small but finite mean free path relative to the gradient scale, $\lambda_{mfp}/L\ll 1$. As illustrated by Fig. 11, however, for many space and astrophysical plasma systems of interest, the scales at which energy is removed from the turbulence (beginning at the small-scale end of the inertial range) are weakly collisional, with $\lambda_{mfp}/L\gtrsim 1$ or $\lambda_{mfp}/L\gg 1$. In these complementary limits relevant to space and astrophysical plasmas, the procedure to calculate the viscosity and resistivity ceases to be valid. In fact, the standard Laplacian viscosity and resistivity terms are likely to be poor approximations to the physical mechanisms that are believed to remove energy from weakly collisional plasma turbulence (see §5.1 for a list of proposed damping mechanisms). Since it is not possible to define clearly the viscosity or resistivity in a weakly collisional system, it is also not possible to define the Reynolds number or magnetic Reynolds number in weakly collisional, kinetic plasma turbulence. Thus, we are forced to choose alternative dimensionless numbers that characterize the plasma and turbulence in weakly collisional systems, so one of the primary aims of this paper is to propose the specific set of fundamental dimensionless Plasma and Turbulence parameters that are suitable for characterizing kinetic plasma turbulence, presented here in Table 3.